Problem 6.05 Mark Harris borrowed some money from his friend and promised to repay him $1,250, $1,340, $1,490, $1,560, and $1,560 over the next five years. If the friend normally discounts investment cash flows at 8.0 percent annually, how much did Mark borrow? (Round answer to 2 decimal places, e.g. 15.25. Do not round factor values.) Present value
Answer #1
Amount borrowed= the present values of the promised repayments at end of Years 1 to 5 --all discounted at 8% p.a.
ie. (1250/1.08^1)+(1340/1.08^2)+(1490/1.08^3)+(1560/1.08^4)+(1560/1.08^5)=
5697.41
Alternately,
1 2 3=1/1.08^n 4=2*3
Year Repayments PV F at 8% PV at 8%
1 1250 0.9259259 1157.41
2 1340 0.8573388 1148.83
3 1490 0.7938322 1182.81
4 1560 0.7350299 1146.65
5 1560 0.6805832 1061.71
PV of amt. Borrowed= 5697.41